OpenStudy (ashy98):

Write the equation of the line that passes through (1, 5) and (−2, 14) in slope-intercept form.

OpenStudy (ashy98):

A.) H = 36w − 4 B.) H = 36w + 4 C.) H = −4w + 36 D.) H = 4w + 36

OpenStudy (sleepyhead314):

slope intercept form can be represented as \( y = mx +b \) where \(m=slope\) and \(b = y~intercept\)

OpenStudy (sleepyhead314):

first, when we are given two points that are on the line, we have to find the \(slope\) do you know how to find slope? :)

OpenStudy (sleepyhead314):

hint: \(\huge slope = \frac{y_2 - y_1}{x_2 - x_1}\) when given the points \((x_1,y_1)\) and \((x_2, y_2)\) so in this case, \(x_1 = 1\) \(y_1 = 5\) \(x_2 = -2\) \(y_2 = 14\)

OpenStudy (sleepyhead314):

Tell me what you get, or if you get stuck so we can go on :)

OpenStudy (ashy98):

@sleepyhead314

OpenStudy (sleepyhead314):

ehhhh? scroll up for a sec, are those the right options? xD